What is the balanced and net ionic equation for the five steps of the copper cycle?
Answer #1

Step 1:

Balanced equation: Cu (s) + 4 HNO3 (aq) = 2 NO2 (aq) + 2 H2O (l) + Cu(NO3)2 (aq)

Net ionic equation: Cu (s) + 4 H+ (aq) + 2 NO3- (aq) = 2 NO2 (aq) + 2 H2O (l) + Cu+2 (aq)

Step 2:

Balanced equation: Cu(NO3)2 (aq) + 2 NaOH (aq) = Cu(OH)2 (aq) + 2 NaNO3 (aq)

Net ionic equation: Cu+2 (aq) + 2 OH- (aq) = Cu(OH)2 (s)

Step 3:

Balanced equation: Cu(OH)2 (s) = CuO (s) + H2O (l)

No net ionic equation

Step 4:

Balanced equation: CuO (s) + H2SO4 (aq) = CuSO4 (aq) + H2O (l)

Net ionic equation: CuO (s) + 2 H+ (aq) = Cu+2 (aq) + H2O (l)

Step 5:

Balanced equation: CuSO4 (aq) + Zn (s) = ZnSO4 (aq) + Cu (s)

Net ionic equation: Cu+2 (aq) + Zn (s) = Zn+2 (aq) + Cu (s)